A certain compound has the following percent composition: C=40.0% and H=6.7%. The empirical formula is...
Answer #1

Ans :-

Element Mass Atomic mass Moles Simplest whole number ratio
C 40 12 40/12 = 3.33 3.33/3.33 = 1
H 6.7 1 6.7/1 = 6.7 6.7/3.33 = 2

Since, the molar ratio of C : H = 1 : 2

Therefore,

The Emperical formula of the compound is = CH2